# NeuS-QA: Grounding Long-Form Video Understanding in Temporal Logic and Neuro-Symbolic Reasoning

### v1

@@ Abstract
Long-Form Video Question Answering (LVQA) poses challenges beyond traditional visual question answering (VQA), which is often limited to static images or short video clips. While current vision-language models (VLMs) perform well in those settings, they struggle with complex queries in LVQA over long videos involving multi-step temporal reasoning and causality. Vanilla approaches, which sample frames uniformly and feed them to a VLM with the question, incur significant token overhead, forcing severe downsampling. As a result, the model often misses fine-grained visual structure, subtle event transitions, or key temporal cues, ultimately leading to incorrect answers. To address these limitations, recent works have explored query-adaptive frame sampling, hierarchical keyframe selection, and agent-based iterative querying. However, these methods remain fundamentally heuristic: they lack explicit temporal representations and cannot enforce or verify logical event relationships. As a result, there are no formal guarantees that the sampled context actually encodes the compositional or causal logic demanded by the question. To address these foundational gaps, we introduce NeuS-QA, a training-free, plug-and-play neuro-symbolic pipeline for LVQA. NeuS-QA translates a natural language question into a formal temporal logic expression, constructs a video automaton from frame-level semantic propositions, and applies model checking to rigorously identify video segments satisfying the question’s logical requirements. Only these logic-verified segments are submitted to the VLM, thus improving interpretability, reducing hallucinations, and enabling compositional reasoning without modifying or fine-tuning the model. Experiments on LongVideoBench and CinePile show NeuS-QA improves performance by over $\mathbf{10\%}$, especially on questions involving event ordering, causality, and multi-step compositional reasoning.
@@ 1 Introduction
Recent advances in video-language models (VLMs) have led to significant progress in visual question answering (VQA) for static images and short video clips (Park et al. 2024; Wang et al. 2024b; Ye et al. 2025; Wang et al. 2025; Choudhury et al. 2023). However, as real-world applications shift toward longer-form content such as surveillance streams, egocentric vlogs, and movie scenes, users are increasingly asking more complex, temporally grounded questions that span multiple events (Wu et al. 2024; Rawal et al. 2024; Choi et al. 2024; Wang et al. 2024a; Tan et al. 2025). For example, as illustrated in Figure 1, a user watching a one-hour travel vlog might ask: “After the man boarded the plane and sat down, what did he grab?”
@@ 2 Related Works
Early VQA systems (Antol et al. 2015) typically relied on fully supervised architectures that encoded entire videos using convolutional and recurrent networks, encoder-decoder pipelines, or attention mechanisms over video features (Zhu et al. 2016; Ye et al. 2017; Jang et al. 2017). However, these models lacked scalability to longer videos and were constrained by limited visual-linguistic reasoning. The emergence of VLMs (Bai et al. 2025; OpenGVLab 2024; Achiam et al. 2023; Li et al. 2024a, b) has enabled zero-shot VQA across diverse domains with stronger generalization. However, these models typically operate by consuming a fixed set of frames sampled from the entire video, limiting their ability to handle long-form videos with complex temporal structures. Their reliance on global frame pooling or simple attention mechanisms often leads to irrelevant or diluted context when presented with multi-event sequences (Wu et al. 2024; Wang et al. 2024a), resulting in degraded performance.
To address the limitations of global frame sampling, recent state-of-the-art (SOTA) approaches adopt retrieval-augmented strategies that identify and condition on the most relevant video segments (Park et al. 2024; Wang et al. 2024b; Ye et al. 2025; Song et al. 2024; Wang et al. 2025; Islam et al. 2025; Choudhury et al. 2023). While effective at reducing irrelevant context, these methods often rely on coarse heuristics or textual summarization, which sacrifice visual fidelity and lack formal structure. In contrast, NeuS-QA uses temporal logic to guide frame selection, yielding a more interpretable and precise grounding of video content.
@@
Symbolic representations have been explored across fields such as robotics (Puranic, Deshmukh, and Nikolaidis 2021; Hasanbeig et al. 2019; Shoukry et al. 2017), autonomous vehicles (Zheng et al. 2025; Mehdipour et al. 2023), generative models (Shah et al. 2025; Sharan et al. 2025; Choi et al. 2025), and structured neural networks (Manginas, Paliouras, and De Raedt 2024; Hashemi et al. 2023). In the context of vision, symbolic methods have been employed to support long-form understanding (Tapaswi et al. 2016; Huang et al. 2020; Xiong et al. 2019; Wu and Krähenbühl 2021); however, these typically rely on latent-space embeddings, sacrificing interpretability. Yang et al. (2023) and Choi et al. (2024) move toward representing videos as formal models to enable structured reasoning. NeuS-QA follows a similar philosophy, modeling videos as verifiable automata to enable interpretable and temporally grounded analysis.
@@ 3 Preliminaries
We represent a video as a stochastic finite-state automaton where each state corresponds to a sampled frame, and transitions capture temporal progression. As videos are a finite sequence of discrete frames with inherent temporal structure, we model them as a discrete-time Markov chain (Norris 1998; Kemeny and Snell 1960).
@@
With formal verification, we assess whether a video automaton satisfies a given TL specification (Clarke, Grumberg, and Peled 1999; Huth and Ryan 2004). We define a path $\pi$ through the automaton $\mathcal{A}_{\mathcal{V}}$ as an infinite sequence of states $\pi=q_{0}q_{1}q_{2}\dots$ such that $\delta(q_{i},q_{i+1})>0$ for all $i\geq 0$. Each path induces a trace—the sequence of label sets assigned to the states along the path, denoted as $\text{trace}(\pi)=\lambda(q_{0})\lambda(q_{1})\lambda(q_{2})\dots$. This trace captures the progression of observable propositions over time, as illustrated in Figure 2. We apply probabilistic model checking (Baier and Katoen 2008) to compute the satisfaction probability $\mathbb{P}[\mathcal{V}\models\varphi]$, quantifying how well the video trace satisfies the TL specification. This provides a formal and interpretable framework for identifying event sequences in video.
@@ 4 Methodology
Vision Foundation Models LLaVA-OneVision-Qwen2-7B (Li et al. 2024a) 32 43.40 64.86 54.39 43.18 53.07
Aria (Li et al. 2024b) 64 52.83 58.11 59.65 36.36 53.07
Qwen2.5-VL-7B-Instruct (Bai et al. 2025) 32 39.62 58.11 54.39 45.46 50.44
GPT-4o-2024-08-06 (Achiam et al. 2023) 32 43.40 58.11 47.37 45.45 49.56
Qwen2-VL-7B-Instruct (Yang et al. 2024) 32 45.28 54.05 45.61 47.73 48.68
InternVL2-8B (OpenGVLab 2024) 32 39.62 48.65 49.12 43.18 45.61
Structured Reasoning Frameworks VideoTree (Wang et al. 2025) 8-32 56.38 50.00 50.00 43.84 50.47
LVNet (Park et al. 2024) 12 41.94 54.17 35.48 46.15 45.59
VideoAgent (Wang et al. 2024b) 32 45.45 52.63 27.27 27.27 40.38
T* (Ye et al. 2025) 8 28.95 35.56 48.33 22.22 35.75
@@
This satisfaction probability is computed using probabilistic model checking, implemented via the STORM model checker (Hensel et al. 2020; Junges and Volk 2021). This process operates under the Probabilistic Computation Tree Logic (PCTL) framework, where $\mu(\cdot)$ analyzes the state labels and transition probabilities in $\mathcal{A}_{\mathcal{V},t}$ to assess how likely the specification $\varphi$ is satisfied.
@@ 5 Experimental Setup
During automaton construction, we use InternVL2-8B (OpenGVLab 2024) as the VLM to generate frame-level detections. The videos are sampled at 3 frames per second. To translate natural language questions into TL specifications, we use the GPT-o1-mini reasoning model (Achiam et al. 2023) with a few-shot prompting strategy and a carefully designed system prompt to capture common temporal structures and patterns. We apply Stormpy (Hensel et al. 2020; Junges and Volk 2021), a probabilistic model checker, to identify video segments that satisfy these TL specifications. We present the results of NeuS-QA as a plug-and-play module with various VLMs in Section 6.
@@
For empirical evaluation of NeuS-QA  we use two LVQA datasets: LongVideoBench (Wu et al. 2024) and CinePile (Rawal et al. 2024). LongVideoBench serves as our primary benchmark. It features videos up to 60 minutes long across diverse domains such as instructional content, egocentric vlogs, and educational streams. CinePile is used as a secondary dataset to demonstrate the performance of NeuS-QA on long-form movie scenes that require reasoning about narrative structure, character interaction, and scene progression. For both datasets, we focus on subsets of question-video pairs that require temporal ordering, causal inference, and compositional logic, filtering out simple examples based solely on static visual recognition. In LongVideoBench, these reasoning-oriented queries correspond to four categories: T3E (event before/after text), E3E (event before/after event), T3O (object before/after text), and O3O (object before/after object). To streamline the evaluation, we burn the provided subtitles into the video, allowing NeuS-QA and consequently any VLM to directly infer and answer questions involving dialogue without requiring separate time-aligned subtitle prompts.
@@
We evaluate NeuS-QA on multiple-choice LVQA tasks using accuracy as the primary metric across both LongVideoBench and CinePile. Since NeuS-QA serves as a plug-and-play key-segment retrieval layer, we apply it modularly on top of several state-of-the-art VLMs, including Qwen2.5 (Bai et al. 2025), GPT-4o (Achiam et al. 2023), and LLaVA-OneVision (Li et al. 2024a). For each model, we report performance with and without our retrieval pipeline to measure the impact of symbolic retrieval. In addition to these baselines, we compare against recent structured LVQA methods such as T*, VideoTree, and LVNet (Ye et al. 2025; Wang et al. 2025; Park et al. 2024), as discussed in Section 2. All benchmarks are conducted using the lmms-eval library (Zhang et al. 2024), which provides a unified framework for evaluating LVQA models across datasets.
@@ 7 Discussion
Another natural question is: “Why temporal logic?” Is it the only or best choice for structuring complex queries? We adopt TL because it strikes a balance between expressivity and compatibility with existing formal verification tools such as Storm and CARL (Hensel et al. 2020; Kremer et al. 2025). TL captures temporal constructs like ordering, simultaneity, and eventuality while remaining amenable to model checking. However, NeuS-QA is logic-agnostic too: alternatives like Linear Temporal Logic (LTL), Signal Temporal Logic (STL), or any domain-specific structured representation could be substituted, as long as they provide a mechanism to create and verify the logical conditions of a query.
@@ 8 Conclusion
We introduce NeuS-QA, a plug-and-play, training-free neuro-symbolic framework for long-form video question answering. By translating natural language queries into temporal logic and verifying them over a video automaton, NeuS-QA isolates segments semantically and temporally aligned with the question before querying a VLM. Our results show strong gains over standard prompting and recent retrieval-based methods (Wang et al. 2025; Park et al. 2024), especially on temporally complex benchmarks like LongVideoBench and CinePile (Wu et al. 2024; Rawal et al. 2024). As VLMs improve, the fidelity of our symbolic alignment will scale in tandem. NeuS-QA is a general framework for structured, interpretable long-video reasoning. The key insight lies not in any particular system design or architecture, but symbolically aligning video content with queries as an effective and principled approach to LVQA.
A.1 Vision Language Model Calibration
We describe the implementation procedures for detecting propositions generated by $\mathcal{L}_{\texttt{Q2TL}}\$and using them to annotate individual frames within synthetic video sequences. The calibration process involves two key components: (1) the inference methodology employed by Vision Language Models (VLMs), and (2) the procedure used to determine optimal calibration thresholds that control the false positive rate for each VLM.
A.2 Vision Language Model Inference Framework
Vision Language Models serve as semantic detection mechanisms. Each atomic proposition $p_{i}\in\mathcal{P}$, including examples such as “person”, “car”, “person in the car”, is processed alongside. Upon receiving binary responses (‘Yes’ or ‘No’) from the VLM, we extract token probabilities from these outputs and utilize them as confidence metrics for detection accuracy.
A.3 Determination of False Positive Rate Thresholds
Calibration Dataset Construction:
The COCO Captions dataset (Chen et al. 2015) serves as the foundation for calibrating several open-source vision language models, including the InternVL2 (OpenGVLab 2024) and Qwen Models (Bai et al. 2025). Since the dataset contains validated image-caption pairs representing positive associations, we generate negative examples by randomly matching images with captions from different image entries within the dataset. The resulting calibration dataset encompasses 40,000 image-caption pairs, each evaluated by the VLM to produce binary ‘Yes’ or ‘No’ classifications.
Optimal Threshold Selection Framework
The identification of optimal VLM thresholds can be approached as either single-class or multi-class classification problems; we implement the multi-class approach. Our methodology begins by aggregating detection results into confidence score arrays paired with one-hot encoded ground truth labels. We systematically evaluate all available confidence scores as potential thresholds, computing prediction accuracy for each candidate threshold. The optimal threshold maximizes overall accuracy, defined as the ratio of combined true positive and true negative predictions to total predictions.
For comprehensive model evaluation, we generate Receiver Operating Characteristic (ROC) curves by calculating true positive rates (TPR) and false positive rates (FPR) across the complete threshold range. Following optimal threshold identification, we apply these calibrated thresholds to VLM predictions.
A.4 Evaluations With Larger Frame Input
We conduct a final evaluation to assess the impact of increasing the number of frames passed to the VLM during answer generation. While NeuS-QA was previously evaluated with 32 input frames (Table 1), we now increase this to 48. As shown in Table 6, using 48 frames consistently improves accuracy across all VLM backbones, further outperforming both the original 32-frame setup and all baseline models. Empirically, 48 frames provides the best performance in our NeuS-QA pipeline.
A.5 Evaluation on All LongVideoBench Categories
We conduct a comprehensive evaluation of NeuS-QA across all 17 LongVideoBench categories, which span a broad range of reasoning skills, including spatial grounding, temporal ordering, object-event interactions, and compositional logic. While our core experiments focused on temporally grounded tasks, this broader analysis tests the generality of NeuS-QA across diverse question structures beyond temporal logic.
As shown in Table 7, NeuS-QA consistently outperforms its respective base vision-language model across nearly all categories for both LLaVA-OneVision and Qwen2.5-VL. The strongest gains appear in temporally complex categories such as T3E, E3E, and T2A, though improvements are also seen in spatial and compositional tasks. These results underscore the versatility of symbolic segment selection, demonstrating its value as a general-purpose reasoning strategy across long-form video question answering.
Method Frames T3E E3E T3O O3O Overall ($\uparrow$)
NeuS-QA + Qwen2.5-VL-7B-Instruct 32 54.72 62.16 66.67 54.55 60.09
NeuS-QA + Aria 32 54.72 64.86 63.16 47.73 58.77
NeuS-QA + GPT-4o-2024-08-06 32 49.06 68.92 59.65 50.00 58.33
NeuS-QA + Qwen2.5-VL-7B-Instruct 48 58.49 62.16 66.67 54.55 60.96
NeuS-QA + Aria 48 54.72 66.22 64.91 52.27 60.53
NeuS-QA + LLaVA-OneVision-Qwen2-7B-ov 48 50.94 77.03 52.63 59.09 61.40
NeuS-QA + GPT-4o-2024-08-06 48 54.72 60.81 56.14 50.00 56.14
Table 6: Effect of Increasing Input Frames. NeuS-QA improves further when increasing from 32 to 48 input frames, showing gains across multiple models on LongVideoBench.
Category LLaVA-OneVision Qwen2.5-VL
Base VLM (%) NeuS-QA (%) Base VLM (%) NeuS-QA (%)
E2O 68.18 70.45 47.73 72.73
E3E 64.86 70.27 52.27 62.16
O2E 56.82 63.64 59.09 65.91
O3O 43.18 52.27 45.46 54.55
S2A 72.73 72.73 63.64 65.91
S2E 65.91 75.00 65.91 70.45
S2O 54.55 61.36 65.91 68.18
SAA 36.36 40.91 43.18 59.09
SOS 59.09 56.82 54.55 63.64
SSS 22.73 20.45 27.27 31.82
T2A 40.91 59.09 47.73 65.91
T2E 50.00 54.55 63.64 68.18
T2O 59.09 65.91 40.91 43.18
T3E 43.40 50.94 39.62 54.72
T3O 54.39 56.14 54.39 66.67
TAA 43.18 47.73 50.00 56.82
TOS 40.91 50.00 45.45 47.73
Overall 51.55 56.96 50.99 59.86
Table 7: Category-wise Accuracy Comparison (%) Across VLMs. NeuS-QA consistently improves performance over its corresponding base VLM across 17 LongVideoBench categories, with especially large gains in temporally complex tasks.

### v2

@@ Abstract
While vision-language models (VLMs) excel at tasks involving single images or short videos, they still struggle with Long Video Question Answering (LVQA) due to its demand for complex multi-step temporal reasoning. Vanilla approaches, which simply sample frames uniformly and feed them to a VLM along with the question, incur significant token overhead. This forces aggressive downsampling of long videos, causing models to miss fine-grained visual structure, subtle event transitions, and key temporal cues. Recent works attempt to overcome these limitations through heuristic approaches; however, they lack explicit mechanisms for encoding temporal relationships and fail to provide any formal guarantees that the sampled context actually encodes the compositional or causal logic required by the question. To address these foundational gaps, we introduce NeuS-QA, a training-free, plug-and-play neuro-symbolic pipeline for LVQA. NeuS-QA first translates a natural language question into a logic specification that models the temporal relationship between frame-level events. Next, we construct a video automaton to model the video’s frame-by-frame event progression, and finally employ model checking to compare the automaton against the specification to identify all video segments that satisfy the question’s logical requirements. Only these logic-verified segments are submitted to the VLM, thus improving interpretability, reducing hallucinations, and enabling compositional reasoning without modifying or fine-tuning the model. Experiments on the LongVideoBench and CinePile LVQA benchmarks show that NeuS-QA significantly improves performance by over $\mathbf{10\%}$, particularly on questions involving event ordering, causality, and multi-step reasoning. We open-source our code at https://utaustin-swarmlab.github.io/NeuS-QA/.
@@ 1 Introduction
Recent advances in video-language models (VLMs) have led to significant progress in visual question answering (VQA) for static images and short video clips (park2024too; wang2024videoagent; ye2025re; wang2025videotree; choudhury2023zero). However, as real-world applications shift toward longer-form content such as surveillance streams, egocentric vlogs, and movie scenes, users are increasingly asking more complex, temporally grounded questions that span multiple events (wu2024longvideobench; rawal2024cinepile; choi2024towards; wang2024lvbench; tan2025allvb). For example, as illustrated in Figure 1, a user watching a one-hour travel vlog might ask: “After the man boarded the plane and sat down, what did he grab?”
@@ 2 Related Works
Early VQA systems (antol2015vqa) typically relied on fully supervised architectures that encoded entire videos using convolutional and recurrent networks, encoder-decoder pipelines, or attention mechanisms over video features (zhu2016visual7w; ye2017video; jang2017tgif). However, these models lacked scalability to longer videos and were constrained by limited visual-linguistic reasoning. The emergence of VLMs (bai2025qwen2; internvl2024v2; achiam2023gpt; li2024llava; li2024aria) has enabled zero-shot VQA across diverse domains with stronger generalization. However, these models typically operate by consuming a fixed set of frames sampled from the entire video, limiting their ability to handle long-form videos with complex temporal structures. Their reliance on global frame pooling or simple attention mechanisms often leads to irrelevant or diluted context when presented with multi-event sequences (wu2024longvideobench; wang2024lvbench), resulting in degraded performance.
To address the limitations of global frame sampling, recent state-of-the-art (SOTA) approaches adopt retrieval-augmented strategies that identify and condition on the most relevant video segments (park2024too; wang2024videoagent; ye2025re; song2024moviechat+; wang2025videotree; islam2025bimba; choudhury2023zero). While effective at reducing irrelevant context, these methods often rely on coarse heuristics or textual summarization, which sacrifice visual fidelity and lack formal structure. In contrast, NeuS-QA uses temporal logic to guide frame selection, yielding a more interpretable and precise grounding of video content.
@@
Symbolic representations have been explored across fields such as robotics (puranic2021learning; hasanbeig2019reinforcement; shoukry2017linear), autonomous vehicles (zheng2025neurostrata; mehdipour2023formal), generative models (shah2025challenge; sharan2025neuro; choi2025we), and structured neural networks (manginas2024nesya; hashemi2023neurosymbolic). In the context of vision, symbolic methods have been employed to support long-form understanding (long1; long2; long3; long-form-video); however, these typically rely on latent-space embeddings, sacrificing interpretability. yang2023specification and choi2024towards move toward representing videos as formal models to enable structured reasoning. NeuS-QA follows a similar philosophy, modeling videos as verifiable automata to enable interpretable and temporally grounded analysis.
@@ 3 Preliminaries
We represent a video as a stochastic finite-state automaton where each state corresponds to a sampled frame, and transitions capture temporal progression. As videos are a finite sequence of discrete frames with inherent temporal structure, we model them as a discrete-time Markov chain (norris1998markov; kemeny1960finite).
@@
With formal verification, we assess whether a video automaton satisfies a given TL specification (clarke1999model; huth2004logic). We define a path $\pi$ through the automaton $\mathcal{A}_{\mathcal{V}}$ as an infinite sequence of states $\pi=q_{0}q_{1}q_{2}\dots$ such that $\delta(q_{i},q_{i+1})>0$ for all $i\geq 0$. Each path induces a trace—the sequence of label sets assigned to the states along the path, denoted as $\text{trace}(\pi)=\lambda(q_{0})\lambda(q_{1})\lambda(q_{2})\dots$. This trace captures the progression of observable propositions over time, as illustrated in Figure 2. We apply probabilistic model checking (Baier2008) to compute the satisfaction probability $\mathbb{P}[\mathcal{V}\models\varphi]$, quantifying how well the video trace satisfies the TL specification. This provides a formal and interpretable framework for identifying event sequences in video.
@@ 4 Methodology
Vision Foundation Models LLaVA-OneVision-Qwen2-7B (li2024llava) 32 43.40 64.86 54.39 43.18 53.07
Aria (li2024aria) 64 52.83 58.11 59.65 36.36 53.07
Qwen2.5-VL-7B-Instruct (bai2025qwen2) 32 39.62 58.11 54.39 45.46 50.44
GPT-4o-2024-08-06 (achiam2023gpt) 32 43.40 58.11 47.37 45.45 49.56
Qwen2-VL-7B-Instruct (yang2024qwen2technicalreport) 32 45.28 54.05 45.61 47.73 48.68
InternVL2-8B (internvl2024v2) 32 39.62 48.65 49.12 43.18 45.61
Structured Reasoning Frameworks VideoTree (wang2025videotree) 8-32 56.38 50.00 50.00 43.84 50.47
LVNet (park2024too) 12 41.94 54.17 35.48 46.15 45.59
VideoAgent (wang2024videoagent) 32 45.45 52.63 27.27 27.27 40.38
T* (ye2025re) 8 28.95 35.56 48.33 22.22 35.75
@@
This satisfaction probability is computed using probabilistic model checking, implemented via the STORM model checker (storm; stormpy). This process operates under the Probabilistic Computation Tree Logic (PCTL) framework, where $\mu(\cdot)$ analyzes the state labels and transition probabilities in $\mathcal{A}_{\mathcal{V},t}$ to assess how likely the specification $\varphi$ is satisfied.
@@ 5 Experimental Setup
During automaton construction, we use InternVL2-8B (internvl2024v2) as the VLM to generate frame-level detections. The videos are sampled at 3 frames per second. To translate natural language questions into TL specifications, we use the GPT-o1-mini reasoning model (achiam2023gpt) with a few-shot prompting strategy and a carefully designed system prompt to capture common temporal structures and patterns. We apply Stormpy (storm; stormpy), a probabilistic model checker, to identify video segments that satisfy these TL specifications. We present the results of NeuS-QA as a plug-and-play module with various VLMs in Section 6.
@@
For empirical evaluation of NeuS-QA  we use two LVQA datasets: LongVideoBench (wu2024longvideobench) and CinePile (rawal2024cinepile). LongVideoBench serves as our primary benchmark. It features videos up to 60 minutes long across diverse domains such as instructional content, egocentric vlogs, and educational streams. CinePile is used as a secondary dataset to demonstrate the performance of NeuS-QA on long-form movie scenes that require reasoning about narrative structure, character interaction, and scene progression. For both datasets, we focus on subsets of question-video pairs that require temporal ordering, causal inference, and compositional logic, filtering out simple examples based solely on static visual recognition. In LongVideoBench, these reasoning-oriented queries correspond to four categories: T3E (event before/after text), E3E (event before/after event), T3O (object before/after text), and O3O (object before/after object). To streamline the evaluation, we burn the provided subtitles into the video, allowing NeuS-QA and consequently any VLM to directly infer and answer questions involving dialogue without requiring separate time-aligned subtitle prompts.
@@
We evaluate NeuS-QA on multiple-choice LVQA tasks using accuracy as the primary metric across both LongVideoBench and CinePile. Since NeuS-QA serves as a plug-and-play key-segment retrieval layer, we apply it modularly on top of several state-of-the-art VLMs, including Qwen2.5 (bai2025qwen2), GPT-4o (achiam2023gpt), and LLaVA-OneVision (li2024llava). For each model, we report performance with and without our retrieval pipeline to measure the impact of symbolic retrieval. In addition to these baselines, we compare against recent structured LVQA methods such as T*, VideoTree, and LVNet (ye2025re; wang2025videotree; park2024too), as discussed in Section 2. All benchmarks are conducted using the lmms-eval library (zhang2024lmms), which provides a unified framework for evaluating LVQA models across datasets.
@@ 7 Discussion
Another natural question is: “Why temporal logic?” Is it the only or best choice for structuring complex queries? We adopt TL because it strikes a balance between expressivity and compatibility with existing formal verification tools such as Storm and CARL (storm; carl). TL captures temporal constructs like ordering, simultaneity, and eventuality while remaining amenable to model checking. However, NeuS-QA is logic-agnostic too: alternatives like Linear Temporal Logic (LTL), Signal Temporal Logic (STL), or any domain-specific structured representation could be substituted, as long as they provide a mechanism to create and verify the logical conditions of a query.
@@ 8 Conclusion
We introduce NeuS-QA, a plug-and-play, training-free neuro-symbolic framework for long-form video question answering. By translating natural language queries into temporal logic and verifying them over a video automaton, NeuS-QA isolates segments semantically and temporally aligned with the question before querying a VLM. Our results show strong gains over standard prompting and recent retrieval-based methods (wang2025videotree; park2024too), especially on temporally complex benchmarks like LongVideoBench and CinePile (wu2024longvideobench; rawal2024cinepile). As VLMs improve, the fidelity of our symbolic alignment will scale in tandem. NeuS-QA is a general framework for structured, interpretable long-video reasoning. The key insight lies not in any particular system design or architecture, but symbolically aligning video content with queries as an effective and principled approach to LVQA.
Acknowledgments
This material is based upon work supported in part by the Office of Naval Research (ONR) under Grant No. N00014-22-1-2254. Additionally, this work was supported by the Defense Advanced Research Projects Agency (DARPA) contract DARPA ANSR: RTX CW2231110. Approved for Public Release, Distribution Unlimited.